# ABACUS: An Electronic Structure Analysis Package for the AI Era

### v1

@@ VI.1 APNS Project
$+\big{[}(\frac{V_{0}}{V})^{2/3}-1\big{]}^{2}\big{[}6-4(\frac{V_{0}}{V})^{2/3}\big{]}\Big{\}}$ (135)
fitted curve. In practical tests, we build three Bravais lattices, including BCC (Body-Centered-Cubic), FCC (Faced-Centered-Cubic ) and diamond for each element, representing different coordination environments, calculate $E$-$V$ datasets and make comparisons with the all-electron results reported in the works of Giovanni P. et al. Wang et al. (2019b). The converged kinetic energy cutoff and $\Delta$ value calculated with it for all available elements in SG15 1.0 pseudopotentialSchlipf and Gygi (2015) are shown in Fig.33.
We perform precision tests on different sets of NAO basis. Similar to PW calculations, the precision is indicated by EOS; additionally, the energy difference of the lowest points of the EOS profile can be an indicator for basis set completeness.
For ABACUS LCAO calculations, the reliability verification test on any given pseudopotential is often slowed by relatively complicated and time-consuming NAOs generation procedure, in which sophisticated experience and insights may be of need, a rapid and highly-automated orbital generation code then becomes necessary. Plus, the precision and efficiency of NAOs are expected to be improved successively, a platform for orbital generation algorithm development with reliable implementation, clear and stable interfaces between modules, high code readability of codes is need. Therefore, an open-source sub-project “ABACUS-ORBGEN” is included as a developing part of APNS. The developing implementation now is still based on the algorithms proposed in Refs. Chen, Guo, and He, 2010a; Lin, Ren, and He, 2021b, but supports orbital configuration automatic setting for pseudopotentials generated with ONCVPSP codes, automatic orbital transferability improving by performing series of bond length scan and potential curve fitting tasks and a quick orbital quality estimation functionality.
VI.2 UniPero Model
The PbTiO3/SrTiO3(PTO/STO) superlattices serve as a model system for exploring real-space topological textures, including flux closures, vortices, skyrmions, and merons. They tested the DPA-1 model by simulating strain-driven topological evolution in the PTO/STO superlattice. Fig. 34(b) illustrates a 40$\times$20$\times$20 supercell, containing 80,000 atoms, employed to model a (PTO)10/(STO)10 superlattice. At a strain state where the in-plane lattice constants $a_{IP}$ = 3.937 Åand $a_{IP}$ = 3.930 Å, the equilibrium state at 300 K obtained with DPMD simulations adopts an ordered polar vortex lattice with alternating vortex and antivortex (Fig. 34(c)). As the in-plane strain increases to $a_{IP}$ = 3.950 Å, the vortex cores shift toward the PTO/STO interfaces (Fig. 34(b)). Finally, at a large tensile in-plane strain ($a_{IP}$ = 3.955 Å), it becomes the periodic electric dipole waves characterized by head-to-tail connected electric dipoles in the form of a sine function (Fig. 34(e)). These results agree with the experimental observations Gong et al. (2021) and previous MD simulations Wu et al. (2023b).
@@ VI.3 DPA-Semi Model
Liu et al. Liu et al. (2024b) generated first-principles data for 19 bulk semiconductors ranging from group IIB to VIA, namely, Si, Ge, SiC, BAs, BN, AlN, AlP, AlAs, InP, InAs, InSb, GaN, GaP, GaAs, CdTe, InTe, CdSe, ZnS, CdS. They used the ABACUS v3.2 package based on the numerical atomic orbitals basis set with the Perdew-Burke-Ernzerhof (PBE) Perdew, Burke, and Ernzerhof (1996) exchange-correlation functional generated atomic datasets to reduce the production cost of the data. The triple-zeta plus double polarization (TZDP) numerical atomic orbitals basis sets were used for all DFT calculations. The atomic datasets are adopted as training data to generate an attention-based deep potential model using the DPA-1 method Zhang et al. (2024c), namely the DPA-Semi model. Fig. 35 shows the procedures for developing the DPA-Semi model.
The total number of data sets generated for 19 semiconductors is around 200k frame. The energy RMSEs of the BAs system are the smallest, which is 0.004 eV/atom, and the force RMSEs of the InSb system are the smallest, which is 0.11 eV/Å. Ref. Liu et al., 2024b exhibits the lattice constants and elastic moduli of 19 semiconductors from group IIB to VIA. The results obtained from the DPA-Semi models are in excellent agreement with the DFT results calculated by ABACUS. These results offer reliable evidence that the DPA-Semi models can be employed to study the physical mechanisms of group IIB to VIA semiconductor systems with ab initio accuracy. The DPA-Semi results of bulk modulus are in excellent agreement with other DFT results Liu et al. (2024b) and available experimental data (Fig. 36). This work provided reliable evidence that the DPA-Semi model can be readily employed to study the scientific issues of group IIB to VIA semiconductor systems with ab initio accuracy.
@@
VI.4 DPA-1 and DPA-2
The Deep Potential team has successively proposed large interatomic potential models like DPA-1 Zhang et al. (2024c) and DPA-2 Zhang et al. (2024d) since 2022. The DPA-1 model is a novel atomic model designed for molecular and materials simulation based on attention mechanism, which is trained on the public OC20 dataset and the materials such as aluminum-magnesium-copper alloys, high entropy alloys, and solid electrolytes. It is applicable to systems containing 56 different elements, and employs a strategy of pretraining combined with fine-tuning to adapt to practical application scenarios.
DPA-2 Zhang et al. (2024d) adopts a multitask training strategy, capable of optimizing datasets from different DFT parameters sources simultaneously. This effectively reduces the model’s requirement for data quality, allowing datasets obtained from different software to be used in the same training process. Moreover, the design of the DPA-2 model enables it to acquire knowledge about chemical composition and configurations during the pretraining phase, which significantly decreases the amount of fine-tuning data needed for specific downstream tasks. In the development of DPA-2, the ABACUS software plays a significant role. It is employed in the DFT calculations for multiple datasets, such as Alloy, FerroEle, SemiCond and high-pressure SuperHydrides.
Alloy
@@ VI.4 DPA-1 and DPA-2
Based on DFT calculations performed using ABACUS, we are developing general LAMs for alloys, covering 53 elements in the periodic table (Li, Na, K, Be, Mg, Ca, Sr, Sc, Y, La, Ti, Zr, Hf, V, Nb, Ta, Cr, Mo, W, Cu, Ag, Au, Zn, Cd, Mn, Re, Fe, Co, Ni, Os, Ir, Pt, Rh, Ru, Pd, Ce, Pr, Nd, Sm, Gd, Tb, Dy, Ho, Er, Tm, Lu, Al, Ga, In, Si, Ge, Sn, Pb (bold ones are commonly used in structural alloys)). The specific pseudopotentials utilized for these elements are listed in lin (2024) and we leveraged the acceleration feature of ABACUS on the Deep Computing Unit (DCU) for all DFT calculations. The DPGEN and APEX workflows generated various structures Zhang et al. (2020c); Li et al. (2024a), including perfect and perturbed crystal structures, vacancies, interstitials, and surfaces of metals and alloys. Some structures were selected through the concurrent learning workflow in DPGEN and then labeled by ABACUS, resulting in a training dataset of $\sim$24,000 entries. By integrating previous training datasets from AIS Square and previous OpenLAM ope (2024a), the latest general LAM for alloys achieved a root mean square error (RMSE) of $\sim$26 meV/atom for energies and $\sim$0.20 eV/Å for atomic forces across the 53 metals and their alloys. This general LAM for alloys demonstrates superior performance in predicting various properties, including lattice parameters, elastic constants, point defects, and surface formation energies, compared to other LAMs Batatia et al. (2024). ABACUS has proven to be a stable and reliable DFT workhorse for diverse structures covering 53 elements and is computationally efficient, particularly on DCU machines.
high-pressure SuperHydrides
@@
During the construction of the dataset for the DPA-1 and DPA-2 models of high-pressure SuperHydrides, all labels were calculated using ABACUS software with plane-wave basis sets. The thermodynamic stability of the discovered structures was confirmed by their energy above the convex hull, which was determined through structure optimization results from ABACUS. Dynamic stability was assessed using the phonon dispersion spectrum calculated by phonopy Togo et al. (2023); Togo (2023), with all forces calculated by ABACUS. Our DFT calculations were performed on DCU clusters, significantly enhancing the efficiency of SCF calculations compared to traditional CPUs. We obtained a dataset comprising 218,349 data frames, covering 29 elements and a pressure range of 150-250 GPa. Utilizing this data, the DPA-1 model achieved a training Root Mean Square Error (RMSE) of 48.1 meV/atom for energy and 334.8 meV/Å for force. The DPA-2 model achieved a training Root Mean Square Error (RMSE) of 55.2 meV/atom for energy and 298.8 meV/Å for force. On the testing dataset, the DPA-1 model yielded an RMSE of 37.6 meV/atom for energy and 171.4 meV/Å for force, and the DPA-2 model yielded an RMSE of 37.4 meV/atom for energy and 122.1 meV/Å for force. Given the large dataset required for the Superhydrides model, ABACUS proved to be the most economical DFT software available. The high-quality training results further attest to the stable performance of ABACUS.
VII Interfaces to Other Packages
VII.1 DeePKS-kit
The DeePKS-kit Chen et al. (2021, 2022a) is designed to train the DeePKS model and can be linked with ABACUS or PySCF Sun et al. (2020). In practice, the training of the DeePKS model is to optimize the following formula
$\rm min_{\omega}\mathbb{E}_{\text{data}}[(E_{\text{label}}-\rm min_{\{\psi_{i}\},\langle\psi_{i}|\psi_{j}\rangle=\delta_{ij}}E_{\text{DeePKS}}[\{\psi_{i}\}|\omega])^{2}].$ (136)
Note that the energy functional $E_{\text{DeePKS}}$ is determined by the parameters $\omega$. When we want to calculate the difference between the label energy ($E_{\text{label}}$) and the ground state DeePKS energy at the specified parameters, we need to obtain it by minimizing the energy functional with respect to wave function $\{\psi_{i}\}$. In other words, every time the parameter is updated, we need to solve a SCF calculation, which is very time-consuming.
To avoid the above problem, we use the projection method to solve Eq. 136. We rewrite it as:
$\min_{\omega}\mathbb{E}_{\text{data}}{[(E_{\text{label}}-E_{\text{DeePKS}}[\{\psi_{i}\}|\omega])^{2}]}$ (137)
$\begin{split}\text{s.t. }&\exists\varepsilon_{i}\leq\mu,\\ &(H[\{\psi_{i}\}|\omega]-\varepsilon_{i})|\psi_{i}\rangle=0,\\ &\langle\psi_{i}|\psi_{j}\rangle=\delta_{ij}\\ &\text{for~{}}\quad i,j=1...N.\end{split}$ (138)
In detail, we optimize the DeePKS model in two iterative steps. The first is the SCF step. Fix model parameters $\omega^{*}$ , and solve Kohn-Sham equation (Eq. 138) to get ground state $\{\psi_{i}^{*}\}$. We usually take $\omega^{*}$ as all zero when the whole training process starts. The second is TRAIN step. Fix orbitals $\{\psi_{i}^{*}\}$, and optimize model parameters $\omega$ following Eq. 137. This way, without the constraints, the optimization for neural network parameters $\omega$ can be carried out quickly. Then, we go back to the first step. We project the orbitals $\{\psi_{i}^{*}\}$ back to the subset that satisfies the constraints by solving the SCF equations. And then update $\omega$. These two steps repeat until convergence is reached.Chen et al. (2021)
We carry out TRAIN step in DeePKS-kit software Chen et al. (2022a), and SCF step in ABACUS. With periodic boundary conditions in ABACUS, we can train and apply DeePKS models in molecular systems and periodic systems. The iterative training process cooperated by DeePKS-kit and ABACUS is summarized in Fig. 37.
@@ VII.1 DeePKS-kit
In the TRAIN step, Eq. 137 corresponds to:
$\min_{\omega}\mathbb{E}_{\text{data}}{[L(\omega)]},$
$L(\omega)=(E_{\text{label}}-E_{\text{DeePKS}}[\{\psi_{i}\}|\omega])^{2}.$ (139)
$L(\omega)$ is the so-called loss function. Loss function $L(\omega)$ can be composed of various energetic terms including the energy $E$, the atomic force $\mathbf{F}$, the stress tensor $\mathbf{\sigma}$ and the band gap $\varepsilon_{g}$ of the interested system, as shown in Fig. 37.
We take atomic force as an example. Force predicted by DeePKS model can be expressed as:
$\bm{F}_{\text{DeePKS}}$ [$[\{\psi_{i}\}|\omega]$ (140)
$=-\frac{\partial E_{\text{DeePKS}}[\{\psi_{i}\}|\omega]}{\partial\mathbf{R}}$
$=\bm{F}_{\text{baseline}}[\{\psi_{i}\}]-\frac{\partial E_{\delta}[\{\psi_{i}\}|\omega]}{\partial\mathbf{R}}$
$=\bm{F}_{\text{baseline}}[\{\psi_{i}\}]-\sum_{Inlmm^{\prime}v}\frac{\partial E_{\delta}}{\partial\mathbf{d}^{I}_{nlv}}\frac{\partial\mathbf{d}^{I}_{nlv}}{\partial D^{I}_{nlmm^{\prime}}}\frac{\partial D^{I}_{nlmm^{\prime}}}{\partial\mathbf{R}}.$
Note that in the TRAIN procedure, we fix orbitals $\{\psi_{i}\}$, so that $\frac{\partial\mathbf{d}^{I}_{nlv}}{\partial\mathbf{X}}=\frac{\partial\mathbf{d}^{I}_{nlv}}{\partial D^{I}_{nlmm^{\prime}}}\frac{\partial D^{I}_{nlmm^{\prime}}}{\partial\mathbf{X}}$ is also fixed, which is outputted in the SCF step. The term $\frac{\partial E_{\delta}}{\partial\mathbf{d}^{I}_{nlv}}$ can be easily calculated in TRAIN step. Once the $\mathbf{F}_{\text{DeePKS}}$ is calculated, the difference between the target force and the current DeePKS force can be optimized.
Similar to the force term, the stress term $\mathbf{\sigma}$ and the band gap term $\varepsilon_{g}$ can be constructed and fed into the loss function. We refer readers to Ref. Ou et al., 2023 for a detailed formula.
VII.2 DeePMD-kit
@@
DeePMD Zhang et al. (2018a, b) is a widely-used Wen et al. (2022) machine-learning molecular dynamics method based on the neural-network potential. ABACUS has an interface with the DeePMD-kit package Wang et al. (2018); Zeng et al. (2023), allowing its deployment as a molecular dynamics engine in executing DeePMD simulations. Trained on system energy, atomic forces, and lattice stress derived from DFT calculations, this neural network potential enables DeePMD to replicate the potential energy surface with first-principle accuracy. Since the neural network potential is computationally much cheaper than the DFT method, DeePMD can simulate large systems at a long time scale, which is barely accessible for AIMD simulations.
@@
Regarding training, DeePMD requires many atomic configurations with system energies, atomic forces, and optionally virial tensors calculated by the DFT software as the training data. Within ABACUS, these requisite data can be obtained either from the self-consistent calculation (section III.2) or from the AIMD simulation (section III.4). An atomic simulation data manipulation Python package dpdatadpd (2024) enables convenient data transfer to the DeePMD training data format.
VII.3 DP-GEN
For more streamlined construction of the Deep Potential PES more efficiently, a recurrent, adaptive learning scheme, DP-GEN (Deep Potential Generator)Zhang et al. (2019, 2020c), is frequently used to generate the training set for DeePMD and train the Deep Potential model. Starting from training with a small initial dataset, DP-GEN iteratively explores unknown configuration space with the trained models and selectively adds new candidate configurations to the training set. This training collection is periodically enriched, ensuring the continued refinement and improvement of the DeePMD models until a designated precision threshold is attained. Such a process indispensably mandates regular DFT-based first-principles computations on the selected candidate configurations, for which integration with ABACUS has been implanted in the DP-GEN package.
VII.4 DeepH

### v3

@@ V.3 DPA-Semi Model
Liu et al. Liu et al. (2024b) generated over 200,000 DFT data for 19 bulk semiconductors ranging from group IIB to VIA, namely, Si, Ge, SiC, BAs, BN, AlN, AlP, AlAs, InP, InAs, InSb, GaN, GaP, GaAs, CdTe, InTe, CdSe, ZnS, CdS. All of the DFT calculations were performed with the ABACUS v3.2 package with the Perdew-Burke-Ernzerhof (PBE) Perdew, Burke, and Ernzerhof (1996) exchange-correlation functional. The triple-zeta plus double polarization (TZDP) NAO basis sets were used for all DFT calculations. The atomic datasets are adopted as training data to generate an attention-based deep potential model using the DPA-1 method Zhang et al. (2024c), namely the DPA-Semi model. Fig. 30 shows the procedures for developing the DPA-Semi model.
To validate the accuracy of the DPA-Semi model, we find the energy RMSEs of the BAs system are the smallest (0.004 eV/atom), while the force RMSEs of the InSb system are the smallest (0.11 eV/Å). For the lattice constants, bulk moduli and elastic constants, the results from DPA-Semi are in excellent agreement with the DFT results, as the bulk moduli data are compared in Fig. 31. This work provided reliable evidence that the DPA-Semi model can be readily employed to study several semiconductor systems with ab initio accuracy.
@@
V.4 DPA-1, DPA-2 and DPA-3
The DPA-1 model Zhang et al. (2024c) is a universal machine learning interatomic potential that leverages the attention mechanism for molecular and materials modeling. Trained on the OC20 benchmark dataset and specialized materials systems, including aluminum-magnesium-copper alloys, high-entropy alloys, and solid-state electrolytes, this architecture achieves robust transferability across diverse atomic environments. With coverage spanning 56 elements in the periodic table, DPA-1 employs a pretraining-fine-tuning paradigm. The attention-based design supports first-principles accuracy while maintaining computational efficiency comparable to classical force fields.
The DPA-2 model Zhang et al. (2024d) employs a multi-task learning framework that effectively integrates datasets from diverse DFT sources. This architecture substantially reduces the model’s dependence on data uniformity, enabling utilization of datasets generated by different methods. Notably, the pretraining phase of DPA-2 incorporates systematic learning of chemical compositions and spatial configurations, which dramatically reduces the demand for task-specific fine-tuning data in downstream applications.
The recently developed DPA-3 Zhang et al. (2025) constructs a hierarchical graph neural network architecture based on line graph transformations, demonstrating substantial improvements in scaling law adherence and cross-dataset generalization capabilities. During the development of both DPA-2 and DPA-3 iterations, the ABACUS software served as a critical computational infrastructure, facilitating DFT calculations across diverse material datasets. Here are two examples.
@@ V.4 DPA-1, DPA-2 and DPA-3
First, based on DFT calculations performed using ABACUS, we are developing general machine-learning-based potential models for alloys, covering 53 elements in the periodic table. The elements are Li, Na, K, Be, Mg, Ca, Sr, Sc, Y, La, Ti, Zr, Hf, V, Nb, Ta, Cr, Mo, W, Cu, Ag, Au, Zn, Cd, Mn, Re, Fe, Co, Ni, Os, Ir, Pt, Rh, Ru, Pd, Ce, Pr, Nd, Sm, Gd, Tb, Dy, Ho, Er, Tm, Lu, Al, Ga, In, Si, Ge, Sn, Pb, where bold ones are commonly used in structural alloys. The specific pseudopotentials utilized for these elements are listed in Ref. lin, 2024 and the calculations were performed efficiently by the ABACUS package on the Deep Computing Unit (DCU) hardware. Specifically, the DP-GEN and APEX workflows generated various structures Zhang et al. (2020c); Li et al. (2025), including perfect and perturbed crystal structures, vacancies, interstitials, and surfaces of metals and alloys. Some structures were selected through the concurrent learning workflow in DP-GEN and then labeled by ABACUS, resulting in a training dataset of $\sim$24,000 entries. By integrating previous training datasets from AI Square and the OpenLAM project ope (2024a), the latest general LAM for alloys achieved a root mean square error (RMSE) of $\sim$26 meV/atom for energies and $\sim$0.20 eV/Å for atomic forces across the 53 metals and their alloys. This general LAM for alloys demonstrates superior performance in predicting various properties, including lattice parameters, elastic constants, point defects, and surface formation energies, compared to other LAMs Batatia et al. (2024). ABACUS has proven to be a stable and reliable DFT workhorse for diverse structures covering these 53 elements.
@@
Second, during the construction of the dataset for the DPA-1 and DPA-2 models of high-pressure SuperHydrides, all labels were calculated using ABACUS with plane-wave basis sets. The thermodynamic stability of the discovered structures was confirmed by their energies above the convex hull, which was determined through structure optimization results from ABACUS. Dynamic stability was assessed using the phonon dispersion spectrum calculated by phonopy Togo et al. (2023); Togo (2023), with all forces calculated by ABACUS. Our DFT calculations were performed on DCU hardware, significantly enhancing the efficiency of SCF calculations compared to traditional CPUs. We obtained a dataset comprising 218,349 data frames, covering 29 elements and a pressure range of 150-250 GPa. Utilizing this data, the DPA-1 model achieved a training RMSE of 48.1 meV/atom for energy and 334.8 meV/Å for force. The DPA-2 model achieved a training RMSE of 55.2 meV/atom for energy and 298.8 meV/Å for force. On the testing dataset, the DPA-1 model yielded an RMSE of 37.6 meV/atom for energy and 171.4 meV/Å for force, while the DPA-2 model yielded an RMSE of 37.4 meV/atom for energy and 122.1 meV/Å for force. Given the large dataset required for the Superhydrides model, ABACUS proved to be the most economical DFT software available. The high-quality training results further attest to the stable performance of ABACUS.
VI Interfaces to Other Packages
VI.1 DeePKS-kit
DeePKS-kit Chen et al. (2021, 2022a) is an innovative computational framework designed to bridge machine learning with quantum-mechanical simulations. The software integrates deep learning techniques with the Kohn-Sham equations to refine exchange-correlation functionals and can be used with ABACUS or PySCF Sun et al. (2020). By training neural networks on high-fidelity quantum chemistry data, DeePKS-kit dynamically corrects errors in conventional DFT approximations, enabling predictions of electronic properties with near-chemical-accuracy at a fraction of the computational cost.
By adopting both DeePKS-kit and ABACUS packages, we optimize the DeePKS model in two iterative steps. As shown in Fig. 32, $L(\omega)$ is the loss function. Loss function $L(\omega)$ can be composed of various energetic terms including the energy $E$, the atomic force $\mathbf{F}$, the stress tensor $\mathbf{\sigma}$ and the band gap $\varepsilon_{g}$ of the interested system The first step is the SCF step, where we fix parameters $\{\omega^{\ast}\}$ in the DeePKS model and solve the Kohn-Sham equation to obtain the ground state electronic wave function. The second step is the TRAIN step, where the electronic wave functions are fixed and the model parameters $\{\omega\}$are optimized. These two steps repeat until convergence is reached. Chen et al. (2021)
@@ VI.1 DeePKS-kit
VI.2 DeePMD-kit
@@
DeePMD Zhang et al. (2018a, b) is a widely-used Wen et al. (2022) machine-learning molecular dynamics method based on the neural-network potential. ABACUS has an interface with the DeePMD-kit package Wang et al. (2018); Zeng et al. (2023), allowing its deployment as a molecular dynamics engine in executing DeePMD simulations. Trained on system energy, atomic forces, and lattice stress derived from DFT calculations, this neural network potential enables DeePMD to replicate the potential energy surface with first-principles accuracy. Since the neural network potential is computationally more efficient than the DFT method, DeePMD can simulate large systems at a long time scale, which is barely accessible for AIMD simulations.
@@
Regarding training, DeePMD requires a large number of atomic configurations with system energies, atomic forces, and optionally virial tensors calculated by the DFT software as the training data. Within ABACUS, these requisite data can be obtained either from the self-consistent calculations (section II.2) or from the AIMD simulations (section II.4). The Python package dpdatadpd (2024) for atomic simulation data manipulation facilitates convenient transfer of data to the DeePMD training data format.
VI.3 DP-GEN
To develop the Deep Potential Energy Surface (PES) more efficiently, the DP-GEN (Deep Potential Generator) framework Zhang et al. (2019, 2020c) employs a recurrent, adaptive learning scheme that systematically refines machine-learned interatomic potentials. The process begins with a minimal initial dataset of atomic configurations labeled via DFT calculations. Through an iterative active learning loop, DP-GEN trains Deep Potential models with varied initializations to explore new regions of the configuration space. During MD simulations, the framework identifies configurations where model predictions diverge significantly and prioritizes these for further DFT validation. This targeted sampling ensures that the training dataset grows strategically, focusing on physically relevant yet under-represented regions of the PES.
The workflow integrates automated DFT computations to validate candidate configurations, DP-GEN has interfaces with first-principles software such as ABACUS, VASP, and Quantum ESPRESSO. As new DFT data points are added, the Deep Potential models undergo retraining, progressively improving their accuracy across diverse atomic environments. The process continues until predefined error thresholds for forces and energies are met. A key innovation of DP-GEN is its parallelized sampling strategy, which enables simultaneous exploration of multiple trajectories to efficiently map high-dimensional configuration spaces while minimizing redundant DFT calculations.
The DP-GEN framework has proven particularly effective for complex systems like multicomponent alloys Wu et al. (2021); Shi et al. (2024), chemical reaction pathways Zeng et al. (2021); Yang et al. (2022), and phase transitions under extreme conditions Gartner et al. (2020); Fu et al. (2023), where traditional sampling methods struggle to capture rare events or subtle energy landscapes. Open-source and modular by design, DP-GEN supports customization for specific materials or molecules, offering a versatile tool for accelerating the development of robust, quantum-mechanically informed potentials.
VI.4 DeepH